Backus, Cooley, & Henriksen | NYU

Capital Inflows: A Threat to Growth?^{∗}

Started: September 28, 2010; Revised: October 6, 2010

Notes for presentation to NYU World Bank conference, “Restoring Growth in Advanced Economies,” October 7, 2010.

Introduction

0. Thank you, I’m delighted to be here.

1. The question for this session is how to restore growth in advanced economies. We tell hundreds of students every year to look at the quality of institutions, and send them to the World Bank’s Doing Business and Governance Indicators to get quantitative measures of them. I think that’s the answer here. We would see, for example, that Greece has weak institutions, something that seems to have surprised some of its creditors.

2. The US and other “advanced” countries have pretty good institutions, on the whole, but it’s clear they’re not perfect. We’ve been reminded, painfully, that financial regulation is inherently difficult to get right, so we’re working on that. The fiscal situation raises red flags in many countries, including the US, and that will have to be dealt with. We can look back in 10-20 years and see how well we did.

3. But I was asked to speak about capital flows. Both here and abroad, the US current account deficit and China’s surplus get a lot of attention. There’s no shortage of hysteria, including that perennial source of economic expertise, the US Congress. The language pushes you this way: Current account “deficits” sound bad, when they get smaller we call it an “improvement” and when they get bigger we call it a “deterioration.” The traditional label, “global imbalances,” suggests that something needs balancing — or that “adjustment” is called for. Like Humpty Dumpty, I’d prefer to control the meaning. I’ll use the more neutral language of international capital flows.

Facts

4. Let’s wipe the slate clean and look at some facts. We know a lot about capital flows: cyclical variation in most countries, political risk in Mexico 1994, oil prices in Saudi Arabia and Norway, and so on.

5. But what are the driving forces behind the historically large net capital flows we see around the world now? Let’s take a look at the evidence:

∗Rough notes, no guarantee of accuracy or sense.

• Current accounts. Ratios to GDP. Substantial dispersion, persistent.

• Net foreign assets. Ditto.

• Spectra. Persistence absolutely clear.

There’s a lot going on here, but it seems clear that movements in net capital flows are long-lasting. The US, for example, has run deficits for most of the last 25 years.

6. The question, then, is what might give rise to persistent movements in capital flows. Business cycles, oil prices, and political crises don’t seem like good candidates: they operate at higher frequency. We need a mechanism that delivers more persistence than this. We think demography is a good candidate. The mechanism, of course, is primarily life-cycle saving. Saving rates in the life cycle model vary systematically by age, so you’d expect to see a link between saving rates and age distributions.

Demography

7. Let’s look at some facts:

• Large differences in dependency rates around the world

• Fertility rates: different, but converging

• Life expectancy: ditto

8. Here’s an example of how that might work. Consider a model in which each country produces the same good with capital and labor. Agents come in overlapping generations, with birth and death rates based on data. Immigration, too. Fixed retirement age. No annuities: Accidental bequests allocated proportionately across surviving agents. Bottom line: strong connection between demography, saving rates, and net capital flows.

Final thoughts

9. So where does that leave us? Are capital flows an “imbalance” to be managed, a reflection of life-cycle saving, or something completely different? We think there’s a role for demographics. Whether you buy that or not, I hope we’ve left you with food for thought.

Random extras

Are capital flows big? NFA? Compared to what?

Test Figure 1 (R fig)

test notes...

Test Figure 2 (Matlab fig)

test notes...

Test Figure 3 (Stata fig)

test notes...